Qn 1. An integer N has digital representation a1a2a3. Moreover,

ˆ None of the digits a1, a2, or a3 and none of the numbers with digital

representation a1a2, a1a3, or a2a3 is divisible by 3.

ˆ N is odd.

ˆ N is divisible by 9.

ˆ a1 ≥ a2 ≥ a3.

Determine all possible numbers N.


1
Expert's answer
2022-02-25T04:25:15-0500

 "N" is odd , and none of the digits "a_1, a_2," or "a_3" is divisible by "3"


"a_3=1, 5, 7"

"N" is divisible by 9


"a_1+a_2+a_3=9k, k=1,2"

Let "a_3=1."

"a_1\ge a_2 \ge a_3," none of the digits "a_1, a_2," or "a_3" is divisible by "3" , and none of the numbers with digital representation "a_1a_2, a_1a_3," or "a_2a_3" is divisible by 3.


"a_2=1, a_3=7"

Or

"a_2=4, a_3=4"

Let "a_3=5."

"a_1\ge a_2 \ge a_3," none of the digits "a_1, a_2," or "a_3" is divisible by "3" , and none of the numbers with digital representation "a_1a_2, a_1a_3," or "a_2a_3" is divisible by 3.


"a_2=5, a_3=8"

Let "a_3=7."

"a_1\ge a_2 \ge a_3," none of the digits "a_1, a_2," or "a_3" is divisible by "3" , and none of the numbers with digital representation "a_1a_2, a_1a_3," or "a_2a_3" is divisible by 3.

It is impossible to find the number.


The possible numbers are "441, 711, 855."